OpenStudy (anonymous):

Which shows the expression below simplified? (8 × 106) × (4 × 10-4) A. 3.2 × 101 B. 3.2 × 104 C. 3.2 × 103 D. 12 × 102

OpenStudy (anonymous):

Using PEMDAS correctly, (8 × 106) × (4 × 10-4) = (848) × (40 - 4) = 848 × 36 = 30528. Using left-to-right sintaxis, (8 × 106) × (4 × 10-4) = (848) × (4 × 6) = 848 × 24 = 20352. A. 3.2 × 101 = 323.2 B. 3.2 × 104 = 332.8 C. 3.2 × 103 = 329.6 D. 12 × 102 = 1224 None of the answers? There must be something wrong with the question. Maybe you missed a sign or a number?